An urn contains 6 white, 4 red and 9 black balls. If 3 balls are drawn at random, find
the probability that: (i) two of the balls drawn are white, (ii) one is of each colour,
(iii) none is red, (iv) at least one is white
1
Expert's answer
2020-03-21T18:11:15-0400

(i) There are 19 balls of which 6 are white, 4 are black and 9 are black.

Number of ways 3 balls can be drawn is "C^3_{19}" .

Number of ways 2 white balls can be drawn is "C^2_6".

Number of ways 1 not white ball can be drawn is "C^1_{13}."

"P(2\; white)=\frac{C^2_6*C^1_{13}}{C^3_{19}}=\frac{6!*13!*3!*16!}{2!*4!*19!}=\frac{65}{323}\approx0.2012."


(ii) Number of ways 1 white ball can be drawn is "C^1_6".

Number of ways 1 red ball can be drawn is "C^1_4" .

Number of ways 1 black ball can be drawn is "C^1_9".

"P(one\; of\; each\; color)=\frac{C^1_6*C^1_4*C^1_9}{C^3_{19}}=\frac{6!*4!*9!*3!*16!}{1!*5!*1!*3!*1!*8!*19!}=\frac{72}{323}\approx0.2229."


(iii) Number of ways 3 not red balls can be drawn is "C^3_{(19-4)}=C^3_{15}"

"P(none\; red)=\frac{C^3_{15}}{C^3_{19}}=\frac{15!*3!*16!}{3!*12!*19!}=\frac{91}{323}\approx0.2817."


(iv) Number of ways 3 not white balls can be drawn is "C^3_{(19-6)}=C^3_{13}" ."P(at\; least\; 1\; white)=1-P(no\; white)=1-\frac{C^3_{13}}{C^3_{19}}=1-\frac{13!*3!*16!}{3!*10!*19!}=1-\frac{286}{969}\approx0.7049."